5(5x+4)-(x-3)=5 solve for x
Answer by sdmmadam@yahoo.com(530)

5(5x+4)-(x-3)=5 ----(1)
5X(5x)+5X(4)-x+3 = 5 (expanding)
25x+20-x+3 = 5
(25x-x)+(20+3) =5
24x+23 = 5
24x= 5-23
24x=-18
x=(-18)/24
x= -(3X6)/(4X6)
x = -3/4
Answer: x =(-3/4)
Verification: LHS of (1) is
5(5x+4)-(x-3)
= 5[5X(-3/4)+4]-[(-3/4)-3]
=5X[(-15/4)+4]+3/4+3
=-75/4+20+3/4+3
=(-75/4+3/4)+20+3
=(-75+3)/4 + 23
=(-72/4)+23
=-18+23
=5
=RHS
Verification seems to be longer than the problem-eh!